HEY! can you help me figure out this digit problem?? Im stuck at the end result and need some help...Thanks!!
The value of a certain two-digit number is 9 times the sum of its digits. If the digits are reversed, the resulting number is 63 less than the original number. Find the original number.

Answer by stanbon(75887)

The value of a certain two-digit number is 9 times the sum of its digits. If the digits are reversed, the resulting number is 63 less than the original number. Find the original number.
----------------
Equations:
10t+u = 9(t+u)
10u+t = (10t+u)-63
-------------------------
Rearrange:
t = 8u
9u-9t = -63
-------------------
t = 8u
u-t = -7
----
Substitute for "t" ans solve for "u":
u-8u = -7
-7u = -7
u = 1
---
Solve for "t":
t = 8*1
t = 8
---
Original Number: 10t+u = 81
===============================
Cheers,
Stan H.
============